OpenStudy (anonymous):

Micah is taking a test. Here is his answer to one of the problems. 3^4 × 3^3 = 9^7 Find and correct his error.

OpenStudy (anonymous):

3^7 You dont multiply out the 3. Justt add the powers in this case

OpenStudy (anonymous):

Hope it makes sense

OpenStudy (anonymous):

u add all of it?

OpenStudy (anonymous):

Im not sure how u got that im confused

OpenStudy (anonymous):

No. So, As an example, 2^3 x 2^2. So, if you simplify we have 8 x 4, correct?

OpenStudy (anonymous):

So we get 32. if we add the exponents and leave the 2, as is, we have then 2^5 ok?

OpenStudy (anonymous):

Multiply it out, 2x2x2x2x2 = 32

OpenStudy (anonymous):

so 2^3 x 2^2 = 2^5 Same goes with 3^4 x 3^3 which is 3^7

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

ohh yeah so how would i write it all out on my homework?

OpenStudy (anonymous):

I would say the answer is reall 3^7 and the mistakes that were made is that the 3's were *'d instead of just being left a 3?

OpenStudy (anonymous):

Micah was not supposed to multiply out the 3. He only needed to add the exponents. The correct answer is 3^4 x 3^3 = 3^7

OpenStudy (anonymous):

Hope this all made sense. Good luck. Message me if you need more help.

OpenStudy (anonymous):

Okay thank u!

OpenStudy (anonymous):

No problem